Parts manufactured by an injection-molding process are subjected to a compressive strength test. We wish to monitor the compressive strength of the parts manufactured by this process, using an x̄ and an s control chart.
Samples of size 9 are to be taken at regular intervals, and their mean compressive strength (in psi = pounds per square inch) and standard deviation are plotted on the charts in time order.
The target values for the compressive strengths are a mean of μ = 90 psi and a standard deviation of σ = 9 psi.

Sample size n c4 2.606 276 2.088 1.964 9515 0.029 1.874 0.9594 0.113 1.806 0.965 0.179 1.751 9693 0.232 1.707 97270.276 1.669 0.7979 8862 0.9213 .94 10

Q1: The center line for the three sigma x̄ control chart would be

  • 9
  • 81
  • 90
  • 99

Q2: The upper control limit for the x̄ control chart would be

  • 117
  • 99
  • 94.5
  • 90

Q3: The lower control limit for the x̄ control chart would be

  • 81
  • 85.5
  • 90
  • 63

Q4: Suppose at the time of sample 10, we observe a mean of 97.5 psi. We should

  • continue sampling, but increase the sample size to 16. The process is barely in control.
  • continue sampling. The process is still in control.
  • declare the process out of control.
  • continue sampling, but reduce the sample size to 4. The process is well in control.

Q5: Suppose at the time of sample 15, we observe a mean of 77.4 psi. We should

  • continue sampling. The process is still in control.
  • continue sampling, but reduce the sample size to 4. The process is well in control.
  • continue sampling, but increase the sample size to 16. The process is barely in control.
  • declare the process out of control.

Q6: Suppose now that we observe 10 sample means in a row that are all below 90. We should

  • continue sampling, but increase the sample size to 16. The process is barely in control.
  • continue sampling. The process is still in control.
  • declare the process out of control.
  • continue sampling, but reduce the sample size to 4. The process is well in control.

Q7: Suppose now that we observe 2 sample means in a row that are all above 90. We should

  • continue sampling. The process is still in control.
  • continue sampling, but reduce the sample size to 4. The process is well in control.
  • continue sampling, but increase the sample size to 16. The process is barely in control.
  • declare the process out of control.

Q8: The center line for the three sigma s chart would be

  • 4.5
  • 18
  • 8.7237
  • 9

Q9: The upper control limit for the three sigma s chart would be

  • 100
  • 4.5
  • 15.363
  • 9

Q10: The lower control limit for the three sigma s chart would be

  • 2.088
  • 18
  • 0
  • 9

I need help figuring out how to answer these questions. I don't just want the answers, I would like to know how to solve for them, as I am feeling completely lost. Please!!
Answer #1

We have ,

μ=90 , sigma = 9 and   n=9

The 3 sigma control limits for X bar chart are given by :

E(X) 35. E. (X)

Rightarrow mu pm 3sigma /sqrt{n}

Center line : CLX-11

Lower Control Limit : LCL_{ar{X}} = mu - 3sigma /sqrt{n}

Upper control Limit :  UCL_{ar{X}} = mu + 3sigma /sqrt{n}

1. Center line = μ=90

2. Upper control Limit = mu + 3sigma /sqrt{n}

= 90 3 9/V9

= 90 + 9

= 99

3. Lower Control Limit   = mu - 3sigma /sqrt{n}

  = 90 -3 9/V9

= 90 - 9

= 81

4. At the time of sample 10, we observe a mean of 97.5 psi. We should "continue sampling, the process is still in control" as it is within control limits.

5. At the time of sample 15, we observe a mean of 77.4 psi. We should "declare the process out of control"

as it is outside the control limits.

6. We observe 10 sample means in a row that are all below 90. We should "continue sampling, but increase the sample size to 16. The process is barely in control." as there is a pattern of 7 or more points below the center line which indicates presence of assignable causes.

7. We observe 2 sample means in a row that are all above 90. We should "continue sampling, the process is still in control." as it is within control limits.

The 3 sigma control limits for three sigma s chart are given by:

E(s) pm 3S.E.(s)

Rightarrow C_{4} .sigma pm 3.C_{5} .sigma

Center line : CL_{s} = C_{4} .sigma

Lower Control Limit : LCL_{s} = C_{4} .sigma - 3.C_{5} .sigma = B_{5}.sigma

Upper control Limit :  UCL_{s} = C_{4} .sigma + 3.C_{5} .sigma = B_{6}.sigma

8. Center line for the three sigma s chart = CL_{s} = C_{4} .sigma

= 0.9693* 9

= 8.7237

9. The upper control limit for the three sigma s chart = UCL_{s} = C_{4} .sigma + 3.C_{5} .sigma = B_{6}.sigma

= 1.707 * 9

= 15.363

10. The lower control limit for the three sigma s chart = LCL_{s} = C_{4} .sigma - 3.C_{5} .sigma = B_{5}.sigma

= 0.232 * 9

= 2.088